OpenStudy (anonymous):

Complete the table and choose the solution for each given value of x in questions 8–13. y = –x2 + 2

OpenStudy (anonymous):

When x = – 1 (1 point) y = –(–1)2 + 1 = 2; (–1, 1) y = (–1)2 + 2 = 1; (–2, 1) y = –(–1)2 + 2 = 1; (–1, 1) y = (–1)2 + 2 = 1; (1, –1)